# GeoReasoner: Geo-localization with Reasoning in Street Views using a Large Vision-Language Model

### v3

@@ 1 Introduction
Street-view geo-localization seeks to predict geographical locations for the given street-view images. The significance of street-view geo-localization is evident in a variety of applications, spanning social studies (Ye et al., 2019b), urban planning (Shen et al., 2018), and navigation (Chalvatzaras et al., 2022). As shown in Figure 1 (left), existing frameworks for street-view geo-localization can be mainly divided into two categories: retrieval-based and classification-based. Retrieval-based approaches entail identifying the most similar image within a geo-tagged image gallery and returning the corresponding geographical location (Zhu et al., 2022; Lin et al., 2022; Zhang et al., 2023b). However, the methods rely on the diversity and comprehensiveness of the geo-tagged image gallery, which can be challenging to curate. Alternatively, classification-based approaches partition the Earth’s surface into distinct regions and assign the input image to a specific region (Clark et al., 2023; Pramanick et al., 2022; Müller-Budack et al., 2018; Seo et al., 2018; Weyand et al., 2016). While these methods leverage shared visual features within a single region, they may neglect valuable semantic information (e.g., signboard texts) crucial for geo-localization. More importantly, these classification methods often operate as black-box models, lacking reasoning capabilities for users to interpret.
@@
Next, we construct an LVLM model, named GeoReasoner, to overcome the difficulty of integrating reasoning capability in geo-localization. The training procedures of GeoReasoner are divided into two folds: reasoning tuning and location tuning. In the first stage, we utilize the 3K reasoned text-image pairs encapsulating human inference knowledge, to fine-tune a well-trained LVLM model with LoRA (Hu et al., 2022) for reasoning adaptation. In the second stage, we leverage the curated 70K high-locatability GSV images dataset, to further fine-tune the LVLM model with another LoRA stacked on the first one for location tuning. We assess GeoReasoner in terms of accuracy for both country-level (i.e., predicting the country in which a street view is located) and city-level (i.e., predicting the city in which a street view is located) geo-localization. The results demonstrate that GeoReasoner outperforms the other counterparts by more than 25% at the country-level geo-localization and 38% at the city-level geo-localization with reasoning on our test dataset. Notably, GeoReasoner performs slightly better than StreetCILP (Haas et al., 2023), which was trained on a substantially larger dataset of 1.1 million geo-tagged street-view images. We also evaluate GeoReasoner against state-of-the-art models for geo-localization using open benchmark datasets. The results show that GeoReasoner achieves comparable performance with only 10k Flickr images used for training. The main contributions of our work are:
@@ 2.1 Street Views
Street views, as the realm of physical environments routinely accessed and engaged with in daily life, bear significant relevance to human perception (Ye et al., 2019b) and urban design (Shen et al., 2018). Analyses of street views contribute to decision-making support (Ye et al., 2019a), improved understanding of urban social and economic structures (Bai et al., 2023b), and traffic asset monitoring and maintenance (Campbell et al., 2019; Li et al., 2021). This study places an emphasis on geo-localization based on street views. Specifically, drawing motivation from Zhang et al. (2018), we delineate the distribution of scene elements to quantify the degree of locatability in street views. Highly locatable street-view images are curated to train an LVLM that surpasses existing geo-localization models.
@@ 2.2 Image-based Geo-localization
Geo-localization entails determining spatial coordinates on the Earth’s surface, with broad applications in practical scenarios, including tracking individual trajectories (Cheng et al., 2022) and positioning autonomous vehicles (Chalvatzaras et al., 2022). This study focuses on image-based geo-localization, utilizing image data as input. Research on image-based geo-localization can be primarily classified into two approaches: retrieval-based (Zhu et al., 2022; Lin et al., 2022; Zhang et al., 2023b) and classification-based (Clark et al., 2023; Pramanick et al., 2022; Müller-Budack et al., 2018; Seo et al., 2018; Weyand et al., 2016).
The retrieval-based approach involves the sequential matching of a single image with a gallery of overhead views, each labeled with geographical coordinates, and identifying the result with the highest matching as the location. However, the utilization of this method is limited due to its requirement for additional reference datasets. The classification-based approach, exemplified by Weyand et al. (2016), involves subdividing the Earth’s surface into thousands of geographical cells and predicting the geographical unit to which an image belongs. The prediction effectiveness can be boosted with a dataset comprising millions of street views, whilst the granularity is influenced by the number of subdivided geographical cells. As such, many studies have been devoted to learning to corresponding multi-level features at different granularity (Vo et al., 2017), or multi-pair features for different tasks (Müller-Budack et al., 2018; Pramanick et al., 2022; Vivanco Cepeda et al., 2024).
We approach image-based geo-localization with a novel paradigm. Specifically, we integrate semantic visual concepts that offer locatable features (Luo et al., 2022; Theiner et al., 2022), and incorporate human reasoning knowledge learned from geo-localization games using an LVLM.
@@ 2.3 Vision-Language Models
The emergence of Large Language Models (LLMs) has significantly impacted various tasks related to natural language processing. These models exhibit remarkable performance in tasks such as text generation (Zhang et al., 2023a) and text-based question answering (Shao et al., 2023), owing to their robust and versatile capabilities. As a result, research attention has shifted towards exploring prompt engineering techniques to enhance the performance of LLMs in downstream tasks (Wei et al., 2022; Yao et al., 2024; Dai et al., 2023; Xu et al., 2023; Ying et al., 2024).
Large vision-language models (LVLMs) integrate visual encoders with LLMs, exhibiting remarkable effectiveness in visual question-answering tasks (Liu et al., 2024; Bai et al., 2023a; Rao et al., 2023). This study harnesses the capabilities of LVLMs to address geo-localization of street views. However, the optimal utilization of LVLMs remains a challenging issue, particularly due to the absence of high-quality training data and a lack of reasoning capabilities. We overcome these challenges through an innovative paradigm and the thoughtful design of model architecture, contributing to a more effective utilization of LVLMs in this domain.
@@ 3.1 Locatability-Enhanced Data Curation
Throughout the development of this work, we observed variations in the degree of locatability among different street views. For example, the images featuring textual signboards or prominent landmarks (e.g., Eiffel Tower) are easily locatable, whilst those captured in a tunnel or obscured by a wall tend to be less locatable. Refer to Figure 4 for further illustration. Simply merging all these street-view images to train an LVLM is not an optimal approach, as the inclusion of poor-quality data can adversely affect the training efficiency of updating an LVLM (Radford et al., 2021). To this end, we introduce locatability, a metric that quantifies the level of locatability of street-view images. We then devise a CLIP-based visual-text pairing network to produce the desired locatability metric for an input street-view image, as shown in Figure 2. The network naturally incorporates data from two perspectives:
@@
  • Textual Clues. Textual clues often serve a pivotal role in delineating the geographical locations of street-view images. Two prominent games, GeoGuessr [^3] https://www.geoguessr.com and Tuxun [^4] https://tuxun.fun, which focus on geo-localization through street views, offer a potential solution to this gap. Their communities have collaboratively curated a well-organized collection of textual clues, used for pinpointing geographical locations across various countries and cities. These clues, maintained by both players and administrators, provide valuable domain knowledge that aids in identifying and evaluating key geographical features in street views. While such datasets now exist (Luo et al., 2022), there are no readily available image-text data pairs specifically tailored for LVLM training. To bridge this gap, we gathered image-text pairs for geo-localization from these two open-source communities. Subsequently, we utilized the BERT-based Named Entity Recognition (NER) (Kenton & Toutanova, 2019) model to clean and filter text that lacked specific geographical location information. In this way, we collected a total of over 3K textual clues that encapsulate rich geo-localization information. For instance, “houses in central Chile are more likely to have terracotta tiled roofs”. Each clue is paired with a corresponding street-view image.
@@
Here, we first use MaskFormer (Cheng et al., 2021) to predict segmentation masks for various classes in GSV images, such as buildings, sky, and vehicles. We then compute an $n$-length vector $\textbf{I}_{seg}$, which quantifies the area ratio of each mask class, where $n$ represents the number of classes. Subsequently, we utilize Sentence-BERT (Reimers & Gurevych, 2019) to measure the similarity between textual clues and semantic segmentation labels, yielding an $m\times n$ matrix $M$, where $m$ is the number of textual clues. After that, we normalize $M$ using min-max normalization, and set values lower than the threshold to zero, resulting in another $m\times n$ matrix $\hat{M}$. We reduce $\hat{M}$ to an $n$-length vector by calculating the mean across its rows, and then normalize it to obtain $\textbf{w}_{loc}$. This vector represents the importance of each semantic segmentation label for geo-localization. With the segmentation mask ratio $\textbf{I}_{seg}$ and the corresponding weight $\textbf{w}_{loc}$, the locatability metric of a GSV image is computed through the multiplication and accumulation of the respective values, as follows:
@@ 3.2 Geo-localization with Reasoning
While many models (e.g.Clark et al. (2023); Pramanick et al. (2022); Müller-Budack et al. (2018); Seo et al. (2018); Weyand et al. (2016)) exist for image-based geo-localization, these models typically predict locations without providing the inference process. This introduces several limitations: First, the models operate as black boxes without providing insights, making it challenging for users to interpret. This obstacle impedes further refinement of the geo-localization model. More importantly, studies have demonstrated that integrating the reasoning process can enhance the capabilities of LLMs (Qiao et al., 2023). Therefore, our objective is to construct an LVLM for image-based geo-localization with reasoning capability.
Model Architecture. Figure 3 illustrates the architecture of the proposed model GeoReasoner, which is based on Qwen-VL (Bai et al., 2023a). GeoReasoner consists of three modules: Vision Encoder, Vision-Language (VL) Adapter and Pre-trained LLM. Specifically, the Vision Encoder module employs the Vision Transformer (ViT) (Dosovitskiy et al., 2021) architecture. The input street-view images are resized to a specific resolution and then divided into a set of image patches. To refine image patches into sequential representations compatible with an LLM, the VL Adapter is introduced. In the VL Adapter, the sequence of visual features is initially condensed to a fixed length to address efficiency challenges posed by the substantial number of visual feature sequences. Subsequently, the processed visual features are integrated with the LLM using cross-attention mechanisms. Following this, the compressed visual feature sequence and text sequence are passed to the Pre-trained LLM module, which functions as a decoder for generating the answer.
@@ 4.1.2 Quantitative Comparison
StreetCLIP (Haas et al., 2023) 0.7943 1.00 0.8854 0.7457 1.00 0.8543
LLaVA (Liu et al., 2024) 0.4029 1.00 0.5744 0.2400 1.00 0.3871
Qwen-VL (Qwen-7B) (Bai et al., 2023a) 0.5829 0.95 0.7225 0.3743 0.89 0.5270
GPT-4V (Achiam et al., 2023) 0.8917 0.34 0.4923 0.5083 0.31 0.3851
ViT* (Dosovitskiy et al., 2021) 0.7100 1.00 0.8304 0.6762 1.00 0.8068
@@ 4.2.1 Qualitative Comparison with SOTA
To assess the efficacy of GeoReasoner in terms of geo-localization with reasoning, we conduct a qualitative comparison with state-of-the-art LVLM-based approaches, including LLaVA (Liu et al., 2024), Qwen-VL (Qwen-7B) (Bai et al., 2023a), and GPT-4V (Achiam et al., 2023). In the experimental phase, we presented the same input street-view images, reasoning process, and result formats to these models. Specifically, a consistent prompt is used, as below:
@@ 4.2.2 Quantitative Comparison with SOTA
We further conduct quantitative experiments to compare with counterparts LVLMs. In addition, we choose StreetCLIP (Haas et al., 2023) as the state-of-the-art classification-based approach and omit retrieval-based approaches relying on a geo-tagged image gallery that is not available. It is important to clarify that, for the LVLM-based approaches, obtaining corresponding and relevant answers is not guaranteed at all times. Therefore, we included Recall rate to measure the proportion of effective answers within the large language models. When calculating the Accuracy rate, only the accuracy of these effective answers is taken into account. We additionally compute F1 values, taking into consideration both Accuracy and Recall metrics.
@@ 4.2.3 Ablation Experiments
To assess the contributions of the location tuning and reasoning tuning components in GeoReasoner, we design several ablation experiments using the Qwen-VL (Bai et al., 2023a) pre-trained model as the baseline. Next, we integrated the Qwen-VL pre-trained model with LoRA1 (GeoReasoner without location tuning) and LoRA2 (GeoReasoner without reasoning tuning). The last experiment involved the full GeoReasoner model, including both the location tuning and reasoning tuning components. The same prompts were utilized for all these models, as in the previous experiments.
@@ 4.2.4 Generalizability Evaluation
To further assess the generalizability of Georeasoner in geo-localization, we conduct additional testing on open Flickr image datasets of Im2GPS (Hays & Efros, 2008) and Im2GPS3k (Vo et al., 2017). Here we use only 10k Flickr images for fine-tuning Georeasoner. Since Georeasoner predicts city names rather than GPS coordinates, we first convert the predicted city names generated by Georeasoner into the GPS coordinates of their respective city centers, then measure the distance between these predicted coordinates and ground-truth locations.
Table 3 and Table 4 present the performance comparison of Georeasoner with PlaNet (Weyand et al., 2016), CPlaNet (Seo et al., 2018), ISNs (Müller-Budack et al., 2018), Translocator (Pramanick et al., 2022), GeoDecoder (Clark et al., 2023), and GeoCLIP (Vivanco Cepeda et al., 2024) on Im2GPS and Im2GPS3k datasets, respectively. The results demonstrate that fine-tuning GeoReasoner using highly locatable images significantly improves prediction accuracy for street, city, and country levels (row 8 vs. row 9 in Table 3, and row 9 vs. row 10 in Table 4). Remarkably, despite being fine-tuned solely on a smaller number of Flickr images, GeoReasoner achieves results comparable to ISNs and GeoCLIP trained on millions of Flickr images, particularly in terms of city- and country-level accuracy. Besides, GeoReasoner trained on the filtered, highly locatable Flickr images also show improvements in the city- and country-level geo-localization, demonstrating the generalizability of our proposed locatability module.

### v4

@@ 1 Introduction
Street-view geo-localization seeks to predict geographical locations for the given street-view images. The significance of street-view geo-localization is evident in a variety of applications, spanning social studies (ye_2019_visual), urban planning (shen_2017_streetvizor), and navigation (chalvatzaras2022). As shown in Figure 1 (left), existing frameworks for street-view geo-localization can be mainly divided into two categories: retrieval-based and classification-based. Retrieval-based approaches entail identifying the most similar image within a geo-tagged image gallery and returning the corresponding geographical location (zhu2022; lin2022; zhang2023). However, the methods rely on the diversity and comprehensiveness of the geo-tagged image gallery, which can be challenging to curate. Alternatively, classification-based approaches partition the Earth’s surface into distinct regions and assign the input image to a specific region (clark2023; pramanick2022; muller2018; seo2018; weyand2016). While these methods leverage shared visual features within a single region, they may neglect valuable semantic information (e.g., signboard texts) crucial for geo-localization. More importantly, these classification methods often operate as black-box models, lacking reasoning capabilities for users to interpret.
@@
Next, we construct an LVLM model, named GeoReasoner, to overcome the difficulty of integrating reasoning capability in geo-localization. The training procedures of GeoReasoner are divided into two folds: reasoning tuning and location tuning. In the first stage, we utilize the 3K reasoned text-image pairs encapsulating human inference knowledge, to fine-tune a well-trained LVLM model with LoRA (hu2021lora) for reasoning adaptation. In the second stage, we leverage the curated 70K high-locatability GSV images dataset, to further fine-tune the LVLM model with another LoRA stacked on the first one for location tuning. We assess GeoReasoner in terms of accuracy for both country-level (i.e., predicting the country in which a street view is located) and city-level (i.e., predicting the city in which a street view is located) geo-localization. The results demonstrate that GeoReasoner outperforms the other counterparts by more than 25% at the country-level geo-localization and 38% at the city-level geo-localization with reasoning on our test dataset. Notably, GeoReasoner performs slightly better than StreetCILP (haas2023learning), which was trained on a substantially larger dataset of 1.1 million geo-tagged street-view images. We also evaluate GeoReasoner against state-of-the-art models for geo-localization using open benchmark datasets. The results show that GeoReasoner achieves comparable performance with only 10k Flickr images used for training. The main contributions of our work are:
@@ 2.1 Street Views
Street views, as the realm of physical environments routinely accessed and engaged with in daily life, bear significant relevance to human perception (ye_2019_visual) and urban design (shen_2017_streetvizor). Analyses of street views contribute to decision-making support (ye2019), improved understanding of urban social and economic structures (bai2023), and traffic asset monitoring and maintenance (campbell2019; li2021caption). This study places an emphasis on geo-localization based on street views. Specifically, drawing motivation from zhang2018, we delineate the distribution of scene elements to quantify the degree of locatability in street views. Highly locatable street-view images are curated to train an LVLM that surpasses existing geo-localization models.
@@ 2.2 Image-based Geo-localization
Geo-localization entails determining spatial coordinates on the Earth’s surface, with broad applications in practical scenarios, including tracking individual trajectories (cheng2022) and positioning autonomous vehicles (chalvatzaras2022). This study focuses on image-based geo-localization, utilizing image data as input. Research on image-based geo-localization can be primarily classified into two approaches: retrieval-based (zhu2022; lin2022; zhang2023) and classification-based (clark2023; pramanick2022; muller2018; seo2018; weyand2016).
The retrieval-based approach involves the sequential matching of a single image with a gallery of overhead views, each labeled with geographical coordinates, and identifying the result with the highest matching as the location. However, the utilization of this method is limited due to its requirement for additional reference datasets. The classification-based approach, exemplified by weyand2016, involves subdividing the Earth’s surface into thousands of geographical cells and predicting the geographical unit to which an image belongs. The prediction effectiveness can be boosted with a dataset comprising millions of street views, whilst the granularity is influenced by the number of subdivided geographical cells. As such, many studies have been devoted to learning to corresponding multi-level features at different granularity (vo2017), or multi-pair features for different tasks (muller2018; pramanick2022; cepeda2023).
We approach image-based geo-localization with a novel paradigm. Specifically, we integrate semantic visual concepts that offer locatable features (luo2022g3; theiner2022), and incorporate human reasoning knowledge learned from geo-localization games using an LVLM.
@@ 2.3 Vision-Language Models
The emergence of Large Language Models (LLMs) has significantly impacted various tasks related to natural language processing. These models exhibit remarkable performance in tasks such as text generation (zhang2023text) and text-based question answering (shao2023), owing to their robust and versatile capabilities. As a result, research attention has shifted towards exploring prompt engineering techniques to enhance the performance of LLMs in downstream tasks (wei2022; yao2023tree; dai2023can; xu2023lvlm; ying2024mmt).
Large vision-language models (LVLMs) integrate visual encoders with LLMs, exhibiting remarkable effectiveness in visual question-answering tasks (liu2023llava; bai2023qwen; rao2023retrieval). This study harnesses the capabilities of LVLMs to address geo-localization of street views. However, the optimal utilization of LVLMs remains a challenging issue, particularly due to the absence of high-quality training data and a lack of reasoning capabilities. We overcome these challenges through an innovative paradigm and the thoughtful design of model architecture, contributing to a more effective utilization of LVLMs in this domain.
@@ 3.1 Locatability-Enhanced Data Curation
Throughout the development of this work, we observed variations in the degree of locatability among different street views. For example, the images featuring textual signboards or prominent landmarks (e.g., Eiffel Tower) are easily locatable, whilst those captured in a tunnel or obscured by a wall tend to be less locatable. Refer to Figure 4 for further illustration. Simply merging all these street-view images to train an LVLM is not an optimal approach, as the inclusion of poor-quality data can adversely affect the training efficiency of updating an LVLM (radford2021learning). To this end, we introduce locatability, a metric that quantifies the level of locatability of street-view images. We then devise a CLIP-based visual-text pairing network to produce the desired locatability metric for an input street-view image, as shown in Figure 2. The network naturally incorporates data from two perspectives:
@@
  • Textual Clues. Textual clues often serve a pivotal role in delineating the geographical locations of street-view images. Two prominent games, GeoGuessr [^3] https://www.geoguessr.com and Tuxun [^4] https://tuxun.fun, which focus on geo-localization through street views, offer a potential solution to this gap. Their communities have collaboratively curated a well-organized collection of textual clues, used for pinpointing geographical locations across various countries and cities. These clues, maintained by both players and administrators, provide valuable domain knowledge that aids in identifying and evaluating key geographical features in street views. While such datasets now exist (luo2022g3), there are no readily available image-text data pairs specifically tailored for LVLM training. To bridge this gap, we gathered image-text pairs for geo-localization from these two open-source communities. Subsequently, we utilized the BERT-based Named Entity Recognition (NER) (kenton2019) model to clean and filter text that lacked specific geographical location information. In this way, we collected a total of over 3K textual clues that encapsulate rich geo-localization information. For instance, “houses in central Chile are more likely to have terracotta tiled roofs”. Each clue is paired with a corresponding street-view image.
@@
Here, we first use MaskFormer (cheng2021) to predict segmentation masks for various classes in GSV images, such as buildings, sky, and vehicles. We then compute an $n$-length vector $\textbf{I}_{seg}$, which quantifies the area ratio of each mask class, where $n$ represents the number of classes. Subsequently, we utilize Sentence-BERT (reimers2019) to measure the similarity between textual clues and semantic segmentation labels, yielding an $m\times n$ matrix $M$, where $m$ is the number of textual clues. After that, we normalize $M$ using min-max normalization, and set values lower than the threshold to zero, resulting in another $m\times n$ matrix $\hat{M}$. We reduce $\hat{M}$ to an $n$-length vector by calculating the mean across its rows, and then normalize it to obtain $\textbf{w}_{loc}$. This vector represents the importance of each semantic segmentation label for geo-localization. With the segmentation mask ratio $\textbf{I}_{seg}$ and the corresponding weight $\textbf{w}_{loc}$, the locatability metric of a GSV image is computed through the multiplication and accumulation of the respective values, as follows:
@@ 3.2 Geo-localization with Reasoning
While many models (e.g.clark2023; pramanick2022; muller2018; seo2018; weyand2016) exist for image-based geo-localization, these models typically predict locations without providing the inference process. This introduces several limitations: First, the models operate as black boxes without providing insights, making it challenging for users to interpret. This obstacle impedes further refinement of the geo-localization model. More importantly, studies have demonstrated that integrating the reasoning process can enhance the capabilities of LLMs (qiao2022reasoning). Therefore, our objective is to construct an LVLM for image-based geo-localization with reasoning capability.
Model Architecture. Figure 3 illustrates the architecture of the proposed model GeoReasoner, which is based on Qwen-VL (bai2023qwen). GeoReasoner consists of three modules: Vision Encoder, Vision-Language (VL) Adapter and Pre-trained LLM. Specifically, the Vision Encoder module employs the Vision Transformer (ViT) (dosovitskiy2020image) architecture. The input street-view images are resized to a specific resolution and then divided into a set of image patches. To refine image patches into sequential representations compatible with an LLM, the VL Adapter is introduced. In the VL Adapter, the sequence of visual features is initially condensed to a fixed length to address efficiency challenges posed by the substantial number of visual feature sequences. Subsequently, the processed visual features are integrated with the LLM using cross-attention mechanisms. Following this, the compressed visual feature sequence and text sequence are passed to the Pre-trained LLM module, which functions as a decoder for generating the answer.
@@ 4.1.2 Quantitative Comparison
StreetCLIP (haas2023learning) 0.7943 1.00 0.8854 0.7457 1.00 0.8543
LLaVA (liu2023llava) 0.4029 1.00 0.5744 0.2400 1.00 0.3871
Qwen-VL (Qwen-7B) (bai2023qwen) 0.5829 0.95 0.7225 0.3743 0.89 0.5270
GPT-4V (gpt4v) 0.8917 0.34 0.4923 0.5083 0.31 0.3851
ViT* (dosovitskiy2020image) 0.7100 1.00 0.8304 0.6762 1.00 0.8068
@@ 4.2.1 Qualitative Comparison with SOTA
To assess the efficacy of GeoReasoner in terms of geo-localization with reasoning, we conduct a qualitative comparison with state-of-the-art LVLM-based approaches, including LLaVA (liu2023llava), Qwen-VL (Qwen-7B) (bai2023qwen), and GPT-4V (gpt4v). In the experimental phase, we presented the same input street-view images, reasoning process, and result formats to these models. Specifically, a consistent prompt is used, as below:
@@ 4.2.2 Quantitative Comparison with SOTA
We further conduct quantitative experiments to compare with counterparts LVLMs. In addition, we choose StreetCLIP (haas2023learning) as the state-of-the-art classification-based approach and omit retrieval-based approaches relying on a geo-tagged image gallery that is not available. It is important to clarify that, for the LVLM-based approaches, obtaining corresponding and relevant answers is not guaranteed at all times. Therefore, we included Recall rate to measure the proportion of effective answers within the large language models. When calculating the Accuracy rate, only the accuracy of these effective answers is taken into account. We additionally compute F1 values, taking into consideration both Accuracy and Recall metrics.
@@ 4.2.3 Ablation Experiments
To assess the contributions of the location tuning and reasoning tuning components in GeoReasoner, we design several ablation experiments using the Qwen-VL (bai2023qwen) pre-trained model as the baseline. Next, we integrated the Qwen-VL pre-trained model with LoRA1 (GeoReasoner without location tuning) and LoRA2 (GeoReasoner without reasoning tuning). The last experiment involved the full GeoReasoner model, including both the location tuning and reasoning tuning components. The same prompts were utilized for all these models, as in the previous experiments.
@@ 4.2.4 Generalizability Evaluation
To further assess the generalizability of Georeasoner in geo-localization, we conduct additional testing on open Flickr image datasets of Im2GPS (hays2008im2gps) and Im2GPS3k (vo2017). Here we use only 10k Flickr images for fine-tuning Georeasoner. Since Georeasoner predicts city names rather than GPS coordinates, we first convert the predicted city names generated by Georeasoner into the GPS coordinates of their respective city centers, then measure the distance between these predicted coordinates and ground-truth locations.
Table 3 and Table 4 present the performance comparison of Georeasoner with PlaNet (weyand2016), CPlaNet (seo2018), ISNs (muller2018), Translocator (pramanick2022), GeoDecoder (clark2023), and GeoCLIP (cepeda2023) on Im2GPS and Im2GPS3k datasets, respectively. The results demonstrate that fine-tuning GeoReasoner using highly locatable images significantly improves prediction accuracy for street, city, and country levels (row 8 vs. row 9 in Table 3, and row 9 vs. row 10 in Table 4). Remarkably, despite being fine-tuned solely on a smaller number of Flickr images, GeoReasoner achieves results comparable to ISNs and GeoCLIP trained on millions of Flickr images, particularly in terms of city- and country-level accuracy. Besides, GeoReasoner trained on the filtered, highly locatable Flickr images also show improvements in the city- and country-level geo-localization, demonstrating the generalizability of our proposed locatability module.